OpenStudy (anonymous):

Help with calculus questions. Answer any of these. Thanks? Let p(x) be a polynomial with integer coefficients and q(x) = x^n * p(x) 1) Suppose p(x) = ax^k . Show that q^(i) (0) is a multiple of n! for any i 2) Suppose p(x) = a_m x^m + ...a1x + a0. Show that q(^i) (0) is a multiple of n! for any i. 3) Use answer in 2 to show that f_n^(i) (0) is an integer for all i.

OpenStudy (anonymous):

this is about the irrationality of pi

OpenStudy (kinggeorge):

For the first one, just combine p(x) and q(x) so that you only have one polynomial. Thus, \[q(x)=x^n ax^k=ax^{n+k}\]Now start taking derivatives.

OpenStudy (kinggeorge):

So \[q'(x)=(n+k)ax^{n+k-1}\]\[q''(x)=(n+k)(n+k-1)ax^{n+k-2}\]In general, \[q^{(i)} ={(n+k)! \over (n+k-i)!}ax^{n+k-i}\]

OpenStudy (kinggeorge):

If we plug in x=0, well, it's always 0. Which is \(0\cdot n!\)

OpenStudy (kinggeorge):

Just do a similar thing for the second problem